OpenStudy (anonymous):

[9.02] What are the solutions of x2 + 8x − 20 = 0? Show all work for full credit.

OpenStudy (anonymous):

x^2 +10x - 2x - 20 = 0 x(x+10) -2 (x+10) = 0 (x-2)(x+10) = 0